OpenStudy (anonymous):

Does anybody know how to find what is an equivalent form of 5+2i/3i? The choices are: a)2-5i/-3 b)-2+5i/3 c)-2-5i/-1 d)2-5i/3 (oh and can you also show how you got the answer please?)

jimthompson5910 (jim_thompson5910):

To rationalize the denominator, you multiply both top and bottom by the conjugate of the denominator. The conjugate of 3i is -3i

jimthompson5910 (jim_thompson5910):

In general, the conjugate of a+bi is a-bi

jimthompson5910 (jim_thompson5910):

Tell me what you get when you do this

OpenStudy (anonymous):

well, 3i and -3i would make -9i^2 which turns into -10

OpenStudy (loser66):

To me, I just time i to both numerator and denominator

OpenStudy (anonymous):

well yes, but when you times it, ohh, ok I think I multiplied wrong

jimthompson5910 (jim_thompson5910):

-9i^2 = -9(-1) = 9 So -9i^2 = 9

OpenStudy (anonymous):

oh yes, um when I mutiplied I got -25i-6i^2/-9i^2, then I simplified that into -15i-7/-10

OpenStudy (anonymous):

oh, that's why, because I added -1 instead of multiplying it

OpenStudy (anonymous):

oh, ok thank-you, again lol

jimthompson5910 (jim_thompson5910):

In the numerator, you will have (5+2i)*(-3i) = -3i(5+2i) (5+2i)*(-3i) = -3i(5)-3i(2i) (5+2i)*(-3i) = -15i - 6i^2 and I'll let you finish